# Discrete Neural Flow Samplers with Locally Equivariant Transformer

### v1

@@ D.4 Locally Equivariant GraphFormer (leGF)
The leGF architecture largely follows the structure of leTF, with the key difference being the computation of attention weights, which are modified to incorporate graph-specific structural biases. Following Ying et al. (2021), given a graph $G=(V,E)$, we define $\psi(v_{i},v_{j})$ as the shortest-path distance between nodes $v_{i}$ and $v_{j}$ if a path exists; otherwise, we assign it a special value (e.g., $-1$). Each possible output of $\psi$ is associated with a learnable scalar $b_{\psi(v_{i},v_{j})}$, which serves as a structural bias term in the self-attention mechanism. Let $A_{i,j}$ denote the $(i,j)$-th element of the Query-Key interaction matrix. The attention weights are then computed as:
@@
where $b_{\psi(v_{i},v_{j})}$ is shared across all attention layers. An illustration of this graph-aware attention mechanism is shown in Figure 10. For further details, we refer the reader to Ying et al. (2021).
@@ E.1.1 Experimental Details
Sampling from Ising Models. We follow the experimental setup described in (Grathwohl et al., 2021, Section F.1). The energy function of the Ising model is given by $\log p(x)\propto ax^{T}Jx+b^{T}x$. In Figure 5, we set $a=0.1$ and $b=0$. The probability path is constructed using a linear schedule with 64 time steps, starting from a uniform initial distribution. The leTF model comprises 3 bidirectional causal attention layers, each with 4 heads and 128 hidden units. Training is performed using the AdamW optimizer with a learning rate of 0.001, a batch size of 128, and for 1,000 epochs (100 steps per epoch). To prevent numerical instability, the log-ratio term is clipped to a maximum value of 5.
Experimental Setup for Figures 1 and 3. In this experiment, we set $a=0.1$ and $b=0.2$. The probability path is defined via a linear schedule over 64 time steps, starting from a uniform initial distribution. The leTF model is composed of 3 bidirectional causal attention layers, each with 4 heads and 64 hidden units. Training is conducted using the AdamW optimiser with a learning rate of 0.001, a batch size of 128, and for 500 epochs (100 steps per epoch). To mitigate numerical instability, the log-ratio term is clipped to a maximum value of 5.
@@ E.1.2 Additional Results
Additional Results of Sampling from Pre-trained EBMs. We present additional results comparing DNFS to baseline methods on sampling from pre-trained EBMs in Figure 11. The results demonstrate that DNFS produces samples that closely resemble those from the oracle distribution.
@@
LEAPS with leConv and leTF. While leConv performs comparably to DNFS on Ising models, its convolutional architecture may struggle to adapt to non-grid data structures. In this experiment, we use the LEAPS [^5] https://github.com/malbergo/leaps algorithm (Holderrieth et al., 2025) to train a neural sampler with two different locally equivariant architectures: leConv (Holderrieth et al., 2025) and leTF (ours). Given the data has 32 dimensions, we pad it to 36 and reshape it into a $6\times 6$ grid to make it compatible with leConv. As shown in Figure 12, LEAPS with leConv fails to achieve meaningful performance, whereas the proposed transformer-based architecture leTF performs comparably to DNFS. This result highlights the limitations of leConv, whose expressiveness is constrained on non-grid data, while leTF offers greater flexibility and generalisation across diverse input structures.
@@
Comparing to MCMC Methods. Compared to MCMC methods, a key advantage of neural samplers is their ability to guarantee convergence when trained optimally, whereas MCMC methods often suffer from slow mixing and poor convergence. To highlight this benefit, we adopt the same setting as in Figure 5, using DNFS with 64 sampling steps. For a fair comparison, we evaluate MCMC baselines with the same number of steps, including Gibbs sampling (Casella & George, 1992), Gradient with Gibbs (GwG) (Grathwohl et al., 2021), and the Discrete Metropolis-adjusted Langevin Algorithm (DMALA) (Zhang et al., 2022b). We present energy histograms based on 5,000 samples in Figure 13, with a long-run Gibbs sampler serving as the oracle. The results show that short-run MCMC methods struggle to produce accurate samples, although gradient-based variants (GwG and DMALA) outperform conventional Gibbs sampling. In contrast, the energy distribution produced by DNFS closely matches the oracle, demonstrating the effectiveness of the proposed neural sampler.
Figure 14: Training loss of different locally equivariant networks (ref: Figure 3).
Figure 15: Training loss comparison between DNFS and LEAPS (ref: Figure 5).
@@
Training Loss Comparison. We train our model by minimising the loss function defined in Equation 5, where a smaller loss value reflects a better fit to the parameterised rate matrix. As such, the loss value serves as a proxy for training quality and model convergence. In Figure 15, we compare the training loss across different locally equivariant network architectures. The results show that the proposed locally equivariant transformer achieves the lowest loss value, indicating its superior capacity to fit the target rate matrix. Furthermore, Figure 15 compares the loss values between DNFS and LEAPS, showing that our approach again outperforms the baseline. This demonstrates the combined effectiveness of our coordinate descent learning algorithm and transformer-based architecture.
@@ E.2.1 Experimental Details
After training, we quantitatively evaluate all methods using negative log-likelihood (NLL) and maximum mean discrepancy (MMD), as reported in Table 2. Specifically, the NLL is computed using the trained EBM on 4,000 samples drawn from the data distribution, with the normalization constant estimated via importance sampling using 1,000,000 samples from a variational Bernoulli distribution with $p=0.5$. For the MMD metric, we follow the protocol in Zhang et al. (2022a), employing an exponential Hamming kernel with a bandwidth of 0.1. All reported results are averaged over 10 independent runs, where each run uses 4,000 samples generated by the trained DNFS.
@@ E.2.2 Additional Results
Additional Results of Probability Mass Estimation. We compare DNFS to various baselines, including PCD (Tieleman, 2008), ALOE+ (Dai et al., 2020), ED-Bern (Schröder et al., 2024), and EB-GFN (Zhang et al., 2022a). As shown in Table 2, DNFS outperforms both Persistent Contrastive Divergence (PCD), which relies on conventional MCMC methods, and the variational approach ALOE+, demonstrating the effectiveness of our method. Additional qualitative results are provided in Figure 16, where DNFS consistently produces accurate energy landscapes and high-quality samples that closely resemble the training data. Furthermore, Figure 17 illustrates the sampling trajectory of DNFS alongside marginal samples from long-run Gibbs sampling. The comparison shows that DNFS produces samples that closely resemble those from Gibbs, demonstrating its ability to approximate the target distribution with high fidelity.
Notably, since the EBM and the sampler are trained jointly, we can evaluate the negative log-likelihood (NLL) of the data using the trained CTMC, as described in Section D.1, and assess sample quality via the maximum mean discrepancy (MMD) using samples generated by the trained sampler. In Table 3, we report results under four evaluation settings: i) NLL (EBM): using the trained EBM with importance sampling; ii) NLL (DNFS): using the trained CTMC following the method in Section D.1; iii) MMD (EBM): samples drawn via Gibbs sampling from the EBM; and iv) MMD (DNFS): samples generated by DNFS. It can be seen that samples generated by DNFS achieved lower MMD compared to those generated by Gibbs sampling, demonstrating the superiority of the learned sampler in capturing the target distribution and producing higher-fidelity samples. However, we observe that DNFS yields less accurate likelihood estimates compared to importance sampling performed with the trained energy function. This is likely because DNFS is trained to satisfy the Kolmogorov equation rather than explicitly optimising the evidence lower bound, as is common in other discrete diffusion models for generative modelling (Shi et al., 2024; Sahoo et al., 2024). Moreover, it is noteworthy that the sample quality and likelihood are not necessarily consistent (Theis et al., 2016, Section 3.2), and DNFS does not directly optimise the likelihood. Thus, the performance of DNFS on NLL is not guaranteed.
@@
Additional Results of Training Ising Models. We further provide a quantitative comparison against baselines for training Ising models. Following Zhang et al. (2022a); Schröder et al. (2024), we evaluate on $D=10\times 10$ grids with $\sigma=0.1,0.2,\dots,0.5$ and $D=9\times 9$ grids with $\sigma=-0.1,-0.2$. Performance is measured by the negative log-RMSE between the estimated $J_{\phi}$ and the true adjacency matrix $J$. As shown in Table 4, while our method underperforms ED-GFN, which is also a neural sampler, it achieves results comparable to Gibbs, GwG, and ED-Bern in most settings, demonstrating its ability to uncover the underlying structure in the data.
@@ E.3.2 Additional Results
DNFS +DLAMA $40.76$ $0.22\%$ 0:08 $93.63$ $0.01\%$ 0:12 $192.30$ $0.92\%$ 0:17
Additional Results on Maximum Cut. We further evaluate DNFS on the maximum cut problem. As shown in Table 5, the trained DNFS significantly outperforms its untrained version, underscoring the effectiveness of our approach. While DNFS slightly lags behind the baselines DMALA and GFlowNet, its MCMC-refined variant achieves the best overall performance, closely approaching the oracle solution provided by Gurobi.
@@
MCMC Refined DNSF. As previously discussed, a key advantage of DNFS is its ability to incorporate additional MCMC steps to refine the sampling trajectory, thanks to the known marginal distribution $p_{t}$. To validate the effectiveness of this MCMC-refined sampling, we conduct an experiment on the MIS problem using the ER16-20 dataset. In this experiment, we compare two methods: DMALA (Zhang et al., 2022b) and DNFS combined with DLAMA, both sampling from the same interpolated distribution $p_{t}\propto p_{0}^{1-t}p_{1}^{t}$. For reference, the average solution size obtained by DNFS without DLAMA refinement is 8.28. As shown in Table 6, applying DLAMA refinement significantly boosts performance, with improvements increasing as more refinement steps are added. More importantly, integrating the proposed neural sampler (DNFS) with the MCMC method (i.e., DNFS + DLAMA) outperforms the standalone MCMC baseline (i.e., DMALA), demonstrating that the learned sampler provides a strong initialisation that guides the refinement process toward better solutions. This result confirms the synergy between neural samplers and MCMC refinement in solving challenging combinatorial problems.
@@
Data 2 Noise $(t:1\rightarrow 0)$ Refer to caption

### v2

@@ D.4 Locally Equivariant GraphFormer (leGF)
The leGF architecture largely follows the structure of leTF, with the key difference being the computation of attention weights, which are modified to incorporate graph-specific structural biases. Following Ying et al. (2021), given a graph $G=(V,E)$, we define $\psi(v_{i},v_{j})$ as the shortest-path distance between nodes $v_{i}$ and $v_{j}$ if a path exists; otherwise, we assign it a special value (e.g., 1-1). Each possible output of $\psi$ is associated with a learnable scalar $b_{\psi(v_{i},v_{j})}$, which serves as a structural bias term in the self-attention mechanism. Let $A_{i,j}$ denote the $(i,j)$-th element of the Query-Key interaction matrix. The attention weights are then computed as:
@@
where $b_{\psi(v_{i},v_{j})}$ is shared across all attention layers. An illustration of this graph-aware attention mechanism is shown in Figure˜10. For further details, we refer the reader to Ying et al. (2021).
@@ E.1.1 Experimental Details
Sampling from Ising Models. We follow the experimental setup described in (Grathwohl et al., 2021, Section F.1). The energy function of the Ising model is given by $\log p(x)\propto ax^{T}Jx+b^{T}x$. In Figure˜5, we set $a=0.1$ and $b=0$. The probability path is constructed using a linear schedule with 64 time steps, starting from a uniform initial distribution. The leTF model comprises 3 bidirectional causal attention layers, each with 4 heads and 128 hidden units. Training is performed using the AdamW optimizer with a learning rate of 0.001, a batch size of 128, and for 1,000 epochs (100 steps per epoch). To prevent numerical instability, the log-ratio term is clipped to a maximum value of 5.
Experimental Setup for Figures˜1 and 3. In this experiment, we set $a=0.1$ and $b=0.2$. The probability path is defined via a linear schedule over 64 time steps, starting from a uniform initial distribution. The leTF model is composed of 3 bidirectional causal attention layers, each with 4 heads and 64 hidden units. Training is conducted using the AdamW optimiser with a learning rate of 0.001, a batch size of 128, and for 500 epochs (100 steps per epoch). To mitigate numerical instability, the log-ratio term is clipped to a maximum value of 5.
@@ E.1.2 Additional Results
Additional Results of Sampling from Pre-trained EBMs. We present additional results comparing DNFS to baseline methods on sampling from pre-trained EBMs in Figure˜11. The results demonstrate that DNFS produces samples that closely resemble those from the oracle distribution.
@@
LEAPS with leConv and leTF. While leConv performs comparably to DNFS on Ising models, its convolutional architecture may struggle to adapt to non-grid data structures. In this experiment, we use the LEAPS [^5] https://github.com/malbergo/leaps algorithm (Holderrieth et al., 2025) to train a neural sampler with two different locally equivariant architectures: leConv (Holderrieth et al., 2025) and leTF (ours). Given the data has 32 dimensions, we pad it to 36 and reshape it into a $6\times 6$ grid to make it compatible with leConv. As shown in Figure˜12, LEAPS with leConv fails to achieve meaningful performance, whereas the proposed transformer-based architecture leTF performs comparably to DNFS. This result highlights the limitations of leConv, whose expressiveness is constrained on non-grid data, while leTF offers greater flexibility and generalisation across diverse input structures.
@@
Comparing to MCMC Methods. Compared to MCMC methods, a key advantage of neural samplers is their ability to guarantee convergence when trained optimally, whereas MCMC methods often suffer from slow mixing and poor convergence. To highlight this benefit, we adopt the same setting as in Figure˜5, using DNFS with 64 sampling steps. For a fair comparison, we evaluate MCMC baselines with the same number of steps, including Gibbs sampling (Casella & George, 1992), Gradient with Gibbs (GwG) (Grathwohl et al., 2021), and the Discrete Metropolis-adjusted Langevin Algorithm (DMALA) (Zhang et al., 2022b). We present energy histograms based on 5,000 samples in Figure˜13, with a long-run Gibbs sampler serving as the oracle. The results show that short-run MCMC methods struggle to produce accurate samples, although gradient-based variants (GwG and DMALA) outperform conventional Gibbs sampling. In contrast, the energy distribution produced by DNFS closely matches the oracle, demonstrating the effectiveness of the proposed neural sampler.
Figure 14: Training loss of different locally equivariant networks (ref: Figure˜3).
Figure 15: Training loss comparison between DNFS and LEAPS (ref: Figure˜5).
@@
Training Loss Comparison. We train our model by minimising the loss function defined in Equation˜5, where a smaller loss value reflects a better fit to the parameterised rate matrix. As such, the loss value serves as a proxy for training quality and model convergence. In Figure˜15, we compare the training loss across different locally equivariant network architectures. The results show that the proposed locally equivariant transformer achieves the lowest loss value, indicating its superior capacity to fit the target rate matrix. Furthermore, Figure˜15 compares the loss values between DNFS and LEAPS, showing that our approach again outperforms the baseline. This demonstrates the combined effectiveness of our coordinate descent learning algorithm and transformer-based architecture.
Table 2: Comparison of the estimated free energy $\mathcal{F}/D$, internel energy $\mathcal{E}/D$, entropy $\mathcal{S}/D$, and effective sample size for Ising models on $D=10\times 10$ grids at different temperatures.
Method ESS Free Energy $\mathcal{F}/D$ Internal Energy $\mathcal{E}/D$ Entropy $\mathcal{S}/D$
$\sigma=0.1$ Optimal Value 11 3.6727-3.6727 0.4282-0.4282 $0.6489$
LEAPS $0.9956\pm 0.0001$ $3.6709\pm 4.9182e^{-5}$ $0.4262\pm 0.0034$ $0.6489\pm 0.0006$
DNFS $0.9985\pm 4.6912e^{-5}$ $3.6709\pm 6.2463e^{-5}$ $0.4271\pm 0.0010$ $0.6488\pm 0.0002$
$\sigma=0.22305$ Optimal Value 11 2.1242-2.1242 1.4763-1.4763 $0.2855$
LEAPS $0.3631\pm 0.1184$ $2.1011\pm 0.0002$ $1.4493\pm 0.0130$ $0.2873\pm 0.0057$
DNFS $0.9685\pm 0.0010$ $2.1120\pm 6.6258e^{-05}$ $1.4743\pm 0.0068$ $0.2811\pm 0.0030$
Quantatitive Results. Recall Equation˜11, where the Ising model is defined as
$p(x)\propto\mathrm{exp}(-E(x))\triangleq\mathrm{exp}(\sigma x^{T}A_{D}x),\quad x\in\{-1,1\}^{D}$ (53)
where $\sigma\in\mathbb{R}$ is the temperature and $A_{D}$ denote the adjacency matrix of the lattice graph. We evaluate our method by comparing the estimated free energy $\mathcal{F}=-\frac{1}{2\sigma}\log Z$, internal energy $\mathcal{E}=\mathbb{E}_{p}[E(x)]$, and entropy $\mathcal{S}=2\sigma(\mathcal{E}-\mathcal{F})$ with their theoretically optimal values derived in Ferdinand & Fisher (1969). The free energy and internal energy are estimated using Sections˜D.1 and 38, respectively. We compare DNFS with LEAPS under two temperature settings, $\sigma=0.1$ and $\sigma=0.22305$, where the latter corresponds to the critical temperature and thus presents a more challenging sampling problem. The result is reported in Table˜2, where we use $2,048$ Monte Carlo samples to estimate the values and report the mean and standard deviation averaged over $10$ independent runs. The results show that DNFS provides accurate estimates of the free and internal energies, closely matching the theoretical value with low variance. In contrast, LEAPS exhibits significantly lower ESS and larger deviations in other metrics under the critical temperature setting, which may indicate insufficient mode coverage. These comparisons highlight the robustness of DNFS and suggest it is less prone to mode collapse across different temperatures.
@@ E.2.1 Experimental Details
After training, we quantitatively evaluate all methods using negative log-likelihood (NLL) and maximum mean discrepancy (MMD), as reported in Table˜3. Specifically, the NLL is computed using the trained EBM on 4,000 samples drawn from the data distribution, with the normalization constant estimated via importance sampling using 1,000,000 samples from a variational Bernoulli distribution with $p=0.5$. For the MMD metric, we follow the protocol in Zhang et al. (2022a), employing an exponential Hamming kernel with a bandwidth of 0.1. All reported results are averaged over 10 independent runs, where each run uses 4,000 samples generated by the trained DNFS.
@@ E.2.2 Additional Results
Additional Results of Probability Mass Estimation. We compare DNFS to various baselines, including PCD (Tieleman, 2008), ALOE+ (Dai et al., 2020), ED-Bern (Schröder et al., 2024), and EB-GFN (Zhang et al., 2022a). As shown in Table˜3, DNFS outperforms both Persistent Contrastive Divergence (PCD), which relies on conventional MCMC methods, and the variational approach ALOE+, demonstrating the effectiveness of our method. Additional qualitative results are provided in Figure˜16, where DNFS consistently produces accurate energy landscapes and high-quality samples that closely resemble the training data. Furthermore, Figure˜17 illustrates the sampling trajectory of DNFS alongside marginal samples from long-run Gibbs sampling. The comparison shows that DNFS produces samples that closely resemble those from Gibbs, demonstrating its ability to approximate the target distribution with high fidelity.
Notably, since the EBM and the sampler are trained jointly, we can evaluate the negative log-likelihood (NLL) of the data using the trained CTMC, as described in Section˜D.1, and assess sample quality via the maximum mean discrepancy (MMD) using samples generated by the trained sampler. In Table˜4, we report results under four evaluation settings: i) NLL (EBM): using the trained EBM with importance sampling; ii) NLL (DNFS): using the trained CTMC following the method in Section˜D.1; iii) MMD (EBM): samples drawn via Gibbs sampling from the EBM; and iv) MMD (DNFS): samples generated by DNFS. It can be seen that samples generated by DNFS achieved lower MMD compared to those generated by Gibbs sampling, demonstrating the superiority of the learned sampler in capturing the target distribution and producing higher-fidelity samples. However, we observe that DNFS yields less accurate likelihood estimates compared to importance sampling performed with the trained energy function. This is likely because DNFS is trained to satisfy the Kolmogorov equation rather than explicitly optimising the evidence lower bound, as is common in other discrete diffusion models for generative modelling (Shi et al., 2024; Sahoo et al., 2024). Moreover, it is noteworthy that the sample quality and likelihood are not necessarily consistent (Theis et al., 2016, Section 3.2), and DNFS does not directly optimise the likelihood. Thus, the performance of DNFS on NLL is not guaranteed.
@@
Additional Results of Training Ising Models. We further provide a quantitative comparison against baselines for training Ising models. Following Zhang et al. (2022a); Schröder et al. (2024), we evaluate on $D=10\times 10$ grids with $\sigma=0.1,0.2,\dots,0.5$ and $D=9\times 9$ grids with $\sigma=-0.1,-0.2$. Performance is measured by the negative log-RMSE between the estimated $J_{\phi}$ and the true adjacency matrix $J$. As shown in Table˜5, while our method underperforms ED-GFN, which is also a neural sampler, it achieves results comparable to Gibbs, GwG, and ED-Bern in most settings, demonstrating its ability to uncover the underlying structure in the data.
@@ E.3.2 Additional Results
DNFS +DMALA $40.76$ $22\% 0:0$8 $93.63$ $0.01\%$ 0:12 $192.30$ $0.92\%$ 0:17
Additional Results on Maximum Cut. We further evaluate DNFS on the maximum cut problem. As shown in Table˜6, the trained DNFS significantly outperforms its untrained version, underscoring the effectiveness of our approach. While DNFS slightly lags behind the baselines DMALA and GFlowNet, its MCMC-refined variant achieves the best overall performance, closely approaching the oracle solution provided by Gurobi.
@@
MCMC Refined DNSF. As previously discussed, a key advantage of DNFS is its ability to incorporate additional MCMC steps to refine the sampling trajectory, thanks to the known marginal distribution $p_{t}$. To validate the effectiveness of this MCMC-refined sampling, we conduct an experiment on the MIS problem using the ER16-20 dataset. In this experiment, we compare two methods: DMALA (Zhang et al., 2022b) and DNFS combined with DLAMA, both sampling from the same interpolated distribution $p_{t}\propto p_{0}^{1-t}p_{1}^{t}$. For reference, the average solution size obtained by DNFS without DLAMA refinement is 8.28. As shown in Table˜7, applying DLAMA refinement significantly boosts performance, with improvements increasing as more refinement steps are added. More importantly, integrating the proposed neural sampler (DNFS) with the MCMC method (i.e., DNFS + DLAMA) outperforms the standalone MCMC baseline (i.e., DMALA), demonstrating that the learned sampler provides a strong initialisation that guides the refinement process toward better solutions. This result confirms the synergy between neural samplers and MCMC refinement in solving challenging combinatorial problems.
Table 8: Results for the maximum independent set problem on RB32-40 graphs with varying parameter pp. Reported values include the solution size (larger is better) and the percentage drop in performance relative to GUROBI (lower is better), indicated in brackets.
$p=$ $0.1$ $0.3$ $0.5$ $0.7$ $0.9$ $1.0$
Gurobi $8.52(0.00\%)$ $8.24(0.00\%)$ $7.06(0.00\%)$ $7.89(0.00\%)$ $8.05(0.00\%)$ $8.63(0.00\%)$
Random $3.38(60.3\%)$ $4.59(44.2\%)$ $4.25(39.8\%)$ $5.54(29.7\%)$ $6.29(21.8\%)$ $6.93(19.6\%)$
DMALA $8.50(0.23\%)$ $8.16(0.97\%)$ $7.01(0.70\%)$ $7.84(0.63\%)$ $8.04(0.12\%)$ $8.63(0.00\%)$
GflowNet $8.20(3.75\%)$ $7.93(3.76\%)$ $6.83(3.25\%)$ $7.78(1.39\%)$ $8.03(0.62\%)$ $8.63(0.00\%)$
DNFS $8.00(6.10\%)$ $7.65(7.16\%)$ $6.67(5.52\%)$ $7.66(2.91\%)$ $7.97(0.99\%)$ $8.63(0.00\%)$
DNFS +DMALA $8.51(0.11\%)$ $8.19(0.60\%)$ $7.02(0.56\%)$ $7.85(0.50\%)$ $8.04(0.12\%)$ $8.63(0.00\%)$
Benchmarking on the RB Graphs. Following Sanokowski et al. (2024), we further evaluate our method on the Maximum Independent Set problem using the RB32-40 graphs, varying the parameter, which controls the problem’s difficulty. Specifically, higher $p$ values of yield easier instances, while lower values result in more challenging graphs. For each setting, we report the solution size, along with the percentage performance drop relative to GUROBI. As shown in Table˜8, the results are consistent with the observation in Table˜1. On easier instances (i.e., higher values of $p$), our method performs competitively, approaching the performance of the oracle solver GUROBI. On more challenging instances, however, DNFS exhibits a performance gap relative to GFlowNet. This gap arises because GFlowNet restricts its sampling trajectories to the feasible solution, effectively narrowing the exploration space. Introducing such an inductive bias into DNFS represents a promising direction for future work. Nevertheless, despite DNFS underperforming GFlowNet in its pure form, it offers a distinct advantage: the intermediate target distribution is explicitly known. This property enables integration with MCMC-based refinement methods (e.g., DNFS+DMALA), which significantly improves performance.
@@
Data to Noise $(t:1\rightarrow 0)$ Refer to caption